# Asymptotically Unbiased Synthetic Control Methods by Distribution Matching

### v3

@@ 5.2 Core Assumptions
We make the following core assumption for identification of the SC weights.
@@
There exists $\widetilde{\Phi}^{*}\in\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$, it holds that
@@
Assumption 5.1 allows us to guarantee the existence of the $G$-th moment of $Y^{N}_{j,t}$ for any $G\in\mathbb{N}$. Assumption 5.2 also has been utilized in existing studies such as Wan et al. (2018), Gunsilius (2023) and Shi et al. (2022). Wan et al. (2018) points out that the SCMs are based on Assumption 5.2. Gunsilius (2023) assumes Assumption 5.2 and estimates $\bm{w}^{*}$ by using minimizing the Wasserstein distance. Shi et al. (2022) investigates the underlying assumptions in the classical SCMs and then find that the combination of fine-grained models and Assumption 5.2 justifies the standard SCMs. We introduce the fine-grained models proposed by Shi et al. (2022) in Section 8.1.
Although the mixture model and linear factor models are different, they are closely related, as pointed out by Shi et al. (2022) and Nazaret et al. (2023). In terms of fine-grained models, Shi et al. (2022) identifies implicit assumptions underlying linear factor models, and we argue that the result justifies the use of a mixture model in the context of linear factor models. Furthermore, Nazaret et al. (2023) independently report this finding. For further details, see Section 8.1.
@@ 5.3 Estimation of SC Weights
Under the assumption of mixture models, we estimate the SC weight to match the characteristic functions (CHFs). We refer to our method as the Density Matching SCM (DMSCM).
Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|.$ (5)
By using the estimate $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}$, we can estimate the treatment effect and distributional treatment effect.
Treatment effect estimation.
Once estimating the weights, as well as the standard SCMs, we can estimate the treatment effect as
@@
We refer to this estimator as the Density Matching SCM (DMSCM) estimator. In Algorithm 1, we summarize our proposed DMSCM estimator with the D2MSCM estimator introduced in the following section.
Distributional treatment effect estimation.
We can also estimate the distributional treatment effect as the counterfactual distribution of $p_{0,t}(y)$ by a bootstrap method. We re-sample $L>1$ observations $\{Y^{l,\dagger}_{0}\}^{L}_{l=1}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$, (i) for each $j\in\mathcal{J}$, we re-sample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$; (ii) we choose $Y^{l,\dagger}_{0}$ from $(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J})$ with probability $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$. In this case, our focus is on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$.
@@ 6 Convergence Analysis
We provide the consistency results about our proposed SC weight estimator. The following theorem describes the results including the consistency for the treatment effect $\tau_{0}$.
@@
Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}^{T_{0}}_{t=1}$ is stationary, strongly mixing, with sum of mixing coefficient bounded by $M$.
@@
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|$
Then, it holds that $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ as $T_{0}\to\infty$. Additionally, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$ and
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right),$
as $T_{0}\to\infty$ and $G\to\infty$.
The proof is shown in Appendix A. Note that this result can be applied for $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case. Thus, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased under each assumption.
@@ 7.1 Conformal Inference on the Treatment Effect
We consider statistical inference on the treatment effect using our proposed DMSCM and D2SCM to test a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Here, we focus on the conformal inference of Chernozhukov et al. (2021), which has been utilized in existing studies, such as Ben-Michael et al. (2021) and Ferman and Pinto (2021).
Our goal is to establish a confidence interval whose $p$-value $\widehat{p}$ satisfies that for any $a\in(0,1)$,
$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a$ (7)
as $T_{0}\to\infty$ and $T_{1}$ is fixed. To achieve this goal, the conformal inference takes the following three steps. First, for the sharp null hypothesis
@@
we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and add it to the original dataset as $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$. Then, we apply our proposed method to the extended datasets consisting of
$\left(Y_{0,1},Y_{0,2},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$ to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and a corresponding predictor of the counterfactual outcome, denoted by $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$. Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pretreatment residuals. Let $u_{t}(\bm{\alpha})$ be the residual defined as $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$ and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$. For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indexes defined as $\pi_{j}(t)=t+j$ if $t+j\leq T$ and $\pi_{j}(t)=t+j-T$ otherwise. Let $\Pi=\{\pi_{1},\dots,\pi_{T-1}\}$. Then, the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
@@
Because $Y^{N}_{0,t}$ is random, by inverting this test, we can construct a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$. The $(1-\xi)$ confidence interval is given as
$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\mid\widehat{p}(\bm{\alpha})>\xi\right\},$
where $\mathcal{A}$ is a set of candidates of $\bm{a}$. Theorem 1 of Chernozhukov et al. (2021) guarantees (7).
@@ 7.2 Two-Sample Homogeneity Test
We define the null and alternative hypotheses as $H_{0}:p^{I}_{t}=p^{N}_{t}$ and $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. Then, we can conduct hypothesis testing by applying two-sample homogeneity test, and our distributional treatment effect estimator. For example, we can employ maximum mean discrepancy (MMD) for testing (Gretton et al., 2012).
@@ 8 Discussion
This section discusses some remaining problems.
@@ 8.1 Justification of Mixture Models
To elucidates assumptions in SCMs, Shi et al. (2022) posits fine-grained models, where each unit has its components, and the outcomes of treated and untreated units are averages of the outcomes of the components. For example, in the empirical study of the tobacco control program in California, the outcome of interest is annual per capita cigarette consumption at the state level, which is an average of the consumption of people in the state. Based on their arguments, we can justify the assumption of mixture models by connecting them to linear factor models.
Assume that there are $U_{j}$ components in unit $j\in\mathcal{J}$. For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ an outcome of a component $u$ in unit $j$ at time $t$. We also suppose an existence of a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$, contributing to individual $u$’s outcome, such as age, education level, or income. We refer to $W_{j,t,u}$ as causes. Assume that $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is an i.i.d. random variable across $u\in\mathcal{U}_{j}$.
The relationship between the causes $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{U}_{j,t})$ can be linear or non-linear, and can also be time-variant. Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, $p_{j,t}(y\mid w)$ be the PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, and $p_{j,t}$ be the marginalized PDF of $W_{j,t}$. The authors assume that interventions are made at a group-level and individuals in each group comply with their group-level intervention.
Under this setting, Shi et al. (2022) finds that the following assumptions are required for the linearity of SCMs (2).
@@
Conditional on the causes $W_{j,t,u}$, the potential outcome $Y^{N}_{j,t,u}$ is independent of the unit $j\in\mathcal{J}$. For population distribution $j\in\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given as $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y|w)$.
@@
Decompose the causes into two subsets $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$. Let $S$ denote the subset that differentiates the treated unit from the selected untreated units (minimal invariant set); that is, its distribution in the target unit is different from its distribution in the selected untreated units. We assume that for all groups, the distribution of $S$ does not change for all time periods $t\leq T$, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
@@
Let $S$ be the minimal invariant set for the target and untreated units. The untreated units are sufficiently similar if the cardinality of the minima invariant set, $\big{|}\mathcal{J}^{U}\big{|}\geq|S|$.
@@
Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there exists at least one untreated unit $j$, where $\zeta_{j}(S=s)>0$.
@@
Suppose that Assumptions 8.18.4 hold. Then, there exists $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{U}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
This proposition also implies that the mixture model holds (Nazaret et al., 2023), which is consistent to the linear factor models. Therefore, based on this result, we can obtain the following result for linear factor models as well as Ferman and Pinto (2021).
@@
Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then, $\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}$ holds as $T_{0}\to\infty$, and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$. Additionally, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converge to $\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$ in probability as $T_{0}\to\infty$ and $G\to\infty$.
The proof is straightforward. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, from Theorem 6.2, the claim holds under (1) with $c_{j}=0$ for $j\in\mathcal{J}$.

### v4

@@ 5.2 Core Assumptions
We now state our core assumption for identifying the SC weights.
@@
There exists $\widetilde{\Phi}^{*}\subset\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$,
@@
Assumption 5.1 ensures that $Y^{N}_{j,t}$ has a finite $G$-th moment for any $G\in\mathbb{N}$. Assumption 5.2 has been employed in earlier work, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Wan et al. (2018) notes that SCMs effectively rely on Assumption 5.2. Gunsilius (2023) similarly assumes (3) and estimates $\bm{w}^{*}$ by minimizing the Wasserstein distance, while Shi et al. (2022) introduces fine-grained models that, combined with Assumption 5.2, can justify standard SCMs. We provide further discussion of these fine-grained models in Section 8.1.
Although mixture models and linear factor models differ, they are closely related (Shi et al., 2022; Nazaret et al., 2023). In the fine-grained models of Shi et al. (2022), certain implicit assumptions underlie linear factor models, suggesting that mixture model assumptions can also be valid for these settings. Independently, Nazaret et al. (2023) arrive at a similar conclusion. We detail these connections in Section 8.1.
@@ 5.3 Estimation of SC Weights
Under the mixture model assumption, we propose estimating SC weights by matching characteristic functions (CHFs). We call our approach the Density Matching SCM (DMSCM).
Let $h\in(0,1)$ be a fixed parameter. We then estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
Using $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect as follows.
5.3.1 Treatment Effect Estimation
Once $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
@@ 5.3.1 Treatment Effect Estimation
We call this estimator the Density Matching SCM (DMSCM) estimator. Algorithm 1 summarizes our proposed estimator and also includes the D2MSCM estimator introduced later.
5.3.2 Distributional Treatment Effect Estimation
To estimate the distributional treatment effect, we can approximate the counterfactual distribution $p_{0,t}(y)$ via bootstrap. Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
  1. 1.
    For each $j\in\mathcal{J}$, resample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$.
  2. 2.
    Choose $Y^{l,\dagger}_{0}$ from $\left(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J}\right)$ with probability $\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$.
Because our primary focus is often on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, this bootstrap procedure enables us to approximate the distribution of the counterfactual outcomes.
@@ 6 Convergence Analysis
We now provide consistency results for our proposed SC weight estimator. The following theorem presents these results, including consistency for the treatment effect $\tau_{0}$.
@@
Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
@@
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
Then, it holds that
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
Additionally, $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$, and
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
The proof is shown in Appendix A. Note that this result also applies to $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case. Consequently, under the respective assumptions, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased.
@@ 7.1 Conformal Inference on the Treatment Effect
We consider statistical inference for the treatment effect using our proposed DMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Specifically, we adopt the conformal inference approach of Chernozhukov et al. (2021), which has been used by, for example, Ben-Michael et al. (2021) and Ferman and Pinto (2021).
Our goal is to construct a confidence interval whose $p$-value $\widehat{p}$ satisfies
$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a\quad\text{for all }a\in(0,1)\quad\text{as }T_{0}\to\infty\text{ and }T_{1}\text{ is fixed}.$ (6)
To achieve this, conformal inference proceeds as follows. First, for the sharp null hypothesis
@@
we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and combine them with the original dataset as
$\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right).$
We then apply our proposed method to this augmented dataset (comprising
$\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$) to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pre-treatment residuals. Define $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$, and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$. For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indices defined by $\pi_{j}(t)=t+j$ if $t+j\leq T$, and $\pi_{j}(t)=t+j-T$ otherwise. Let $\Pi=\{\pi_{1},\dots,\pi_{T-1}\}$. Then the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
@@
Because $Y^{N}_{0,t}$ is random, inverting this test yields a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$. The $(1-\xi)$ confidence interval is
$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\bigm{|}\widehat{p}(\bm{\alpha})>\xi\right\},$
where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6).
@@ 7.2 Two-Sample Homogeneity Test
To test distributional treatment effects, consider the null and alternative hypotheses $H_{0}:p^{I}_{t}=p^{N}_{t}$ versus $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. One can perform a two-sample homogeneity test in combination with our distributional treatment effect estimator. For instance, one may employ maximum mean discrepancy (MMD) (Gretton et al., 2012) to implement this hypothesis test.
@@ 8 Discussion
We conclude by discussing some remaining issues.
@@ 8.1 Justification of Mixture Models
To clarify assumptions in SCMs, Shi et al. (2022) proposes fine-grained models wherein each unit is composed of multiple components, and the outcomes of treated and untreated units are averages of the outcomes across these components. For example, in the empirical study of California’s tobacco control program, the outcome of interest (annual per capita cigarette consumption) at the state level is an average of individuals’ consumption in the state. Building on this framework, one can justify mixture model assumptions by relating them to linear factor models.
Assume there are $U_{j}$ components in unit $j\in\mathcal{J}$. For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ be the potential outcomes of component $u$ in unit $j$ at time $t$. Suppose there is a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$ capturing individual characteristics (e.g., age or income). We assume $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is i.i.d. across $u\in\mathcal{U}_{j}$.
The relationship between $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ can be linear or nonlinear, and possibly time-varying. Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, and let $p_{j,t}(y\mid w)$ be the conditional PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, while $p_{j,t}$ is the marginal PDF of $W_{j,t}$. We assume that the intervention is administered at a group level and that individuals within each group comply with their group-level treatment.
Under these conditions, Shi et al. (2022) shows that the following assumptions are required for the linearity of SCMs in (2):
@@
Conditional on the causes $W_{j,t,u}$, the potential outcome $Y^{N}_{j,t,u}$ is independent of the unit $j\in\mathcal{J}$. For the population distribution $j\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given by $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y\mid w)$.
@@
Decompose the causes as $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$. Let $S$ denote the subset of causes that differentiates the treated unit from the selected untreated units (the minimal invariant set). We assume that, for all groups, the distribution of $S$ does not change for all time periods $t\leq T$. Hence, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
@@
Let $S$ be the minimal invariant set for the target and untreated units. The untreated units are sufficiently similar if the cardinality of this minimal invariant set satisfies $\left|\mathcal{J}^{U}\right|\geq\lvert S\rvert$.
@@
Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there is at least one untreated unit $j$ with $\zeta_{j}(S=s)>0$.
@@
Suppose Assumptions 8.18.4 hold. Then there exist weights $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
This result also implies that the mixture model (3) holds (Nazaret et al., 2023), aligning it with linear factor models. Hence, we obtain the following corollary for linear factor models similar to Ferman and Pinto (2021):
@@
Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then
$\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converges to
$\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$
in probability as $T_{0}\to\infty$ and $G\to\infty$.
Proof Sketch. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, applying Theorem 6.2 yields the desired result under (1) with $c_{j}=0$ for each $j\in\mathcal{J}$.